# arXiv:2603.00316v1[physics.flu-dyn] 27 Feb 2026

### v1

@@ 3.4 Evolution of unstable planar flames in the presence of a uniform flow
$b$ $=\frac{\mathfrak{r}-1}{1+\mathfrak{m}}\mathcal{M}_{t}+\frac{\mathfrak{r}+\mathfrak{m}}{1+\mathfrak{m}}\mathcal{M}_{c}+\left(\frac{\mathfrak{r}+1}{1+\mathfrak{m}}\mathcal{M}_{t}-\frac{\mathfrak{r}+\mathfrak{m}}{1+\mathfrak{m}}\mathcal{M}_{c}\right)\mathcal{V}$
$\quad+\left(\frac{\mathfrak{r}-1}{1+\mathfrak{m}}\frac{\mathfrak{m}}{\mathfrak{r}}\mathcal{M}_{t}+\frac{\mathfrak{r}+\mathfrak{m}}{1+\mathfrak{m}}\mathcal{M}_{g}\right)\mathrm{g},$ (38)
@@
$b$ $=\frac{\mathfrak{r}-1}{1+\mathfrak{m}}\left(1+\frac{\mathfrak{m}\mathrm{g}}{\mathfrak{r}}\right)\mathcal{M}_{t}+\frac{\mathfrak{r}+\mathfrak{m}}{1+\mathfrak{m}}(\mathcal{M}_{c}+\mathcal{M}_{g}\mathrm{g}).$ (41)
@@
$b$ $=\frac{2\mathfrak{r}\mathcal{M}_{t}}{1+\mathfrak{m}}+\left(\frac{\mathfrak{r}-1}{1+\mathfrak{m}}\frac{\mathfrak{m}\mathcal{M}_{t}}{\mathfrak{r}}+\frac{\mathfrak{r}+\mathfrak{m}}{1+\mathfrak{m}}\mathcal{M}_{g}\right)\mathrm{g}.$ (43)
A key observation is that the expressions for $\mathcal{V}=1$ no longer involves the curvature Markstein number $\mathcal{M}_{c}$ as defined in formula (10). This is because the factor $(1-\mathbf{v}\cdot\mathbf{n})\nabla\cdot\mathbf{n}$ vanishes then to linear order in the stability analysis, leaving only $\mathcal{M}_{t}$ and $\mathcal{M}_{g}$ contributions. Another aspect worth emphasising is that viscosity variations ($\mu\neq 1$) are ultimately responsible for the dependence of the dispersion relation on the uniform flow magnitude $\mathcal{V}$; indeed, when $\mu=1$, we have $\mathcal{M}_{c}=\mathcal{M}_{t}$ and $\mathcal{V}$ disappears from equations (36) to (39). Consequently, when $\mu=1$, the linear stability of the flame becomes insensitive to the presence of a uniform flow, recovering the usual behaviour common to unconfined flames.
@@ 3.5 Weakly nonlinear regime: the Michelson–Sivashinsky equation for Darcy flows
Near the instability threshold ($a\gtrsim 0$), the unstable wavenumbers scale as $k=O(a)$ when $a\to 0^{+}$, and the growth rate becomes $s\approx a|k|-bk^{2}=O(a^{2})$, with $b=\mathcal{M}_{c}(1-\mathcal{V})+\mathcal{V}\mathcal{M}_{t}/(1+\mathfrak{m})$ remaining $O(1)$. This scaling mirrors that of the classical Darrieus–Landau instability, but the coefficients now incorporate Darcy-specific contributions. Following the approach of Sivashinsky ashinsky1988nonlinear, michelson1977nonlinear, one obtains a Michelson–Sivashinsky (MS) equation for the flame shape deviation $f(x,t)$ from its unperturbed planar value $y=(\mathcal{V}-1)t$, namely
@@ 4 Flame stability based on Euler–Darcy’s description
$c_{1}$ $=2k+[\mathcal{M}_{c}^{\prime}+(\mathfrak{r}-1)\mathcal{M}_{t}]2k^{2}+\varphi(1+ck),$ (47)
$c_{0}$ $=-a\varphi k+\big(b\varphi+1-\mathfrak{r}\big)k^{2}+Dk^{3}$ (48)
@@
$\mathcal{M}_{c}^{\prime}$ $=\mathcal{M}_{c}(1-\mathcal{V})+\mathcal{V}\mathcal{M}_{t},\quad\varphi=\frac{12\textit{Pr}(1+\mathfrak{m})}{\varepsilon^{2}\mathfrak{m}},$ (49)
$D$ $=2\mathfrak{r}\mathcal{M}_{c}^{\prime}+(\mathfrak{r}-1)\mathcal{M}_{t},\quad\varepsilon=\frac{\sqrt{12\kappa_{u}}}{\delta_{L}}.$ (50)
@@
The onset criterion for instability of the planar flame is now given as a function of $\varepsilon$ (or $\varphi$). Following daou2025hydrodynamic, we assume $\mathcal{M}_{c}^{\prime},\mathcal{M}_{t}>0$ (so that $c_{1},c_{2},D>0$). The flame is unstable when $a>a_{c}$, with
$a_{c}$ $=\begin{cases}0&\text{if }\varphi>\varphi_{c},\\ \displaystyle\frac{-b(\mathfrak{r}-1)}{4D}\,\frac{\varphi_{c}}{\varphi}\left(1-\frac{\varphi}{\varphi_{c}}\right)^{\!2}&\text{if }\varphi<\varphi_{c}\end{cases}$ (51)
@@
$k_{c}$ $=\begin{cases}0&\text{if }\varphi>\varphi_{c},\\ \displaystyle\frac{\mathfrak{r}-1}{2D}\left(1-\frac{\varphi}{\varphi_{c}}\right)&\text{if }\varphi<\varphi_{c}.\end{cases}$ (52)
The first criterion, $a>a_{c}=0$, applies under strong confinement (large $\varphi$) and is predicted by Darcy’s law; it remains valid up to the critical value $\varphi_{c}$. For $\varphi<\varphi_{c}$, the wavenumber at onset is non-zero, with $k_{c}\to(\mathfrak{r}-1)/2D$ as $\varphi/\varphi_{c}\to 0$, corresponding to unconfined flames.
@@
where $\alpha=\varphi k_{c}/c_{1}(k_{c})$ and $\gamma=Dk_{c}/c_{1}(k_{c})$ are positive numbers. For $\varphi>\varphi_{c}$, the flame dynamics is governed by a Michelson–Sivashinsky-type equation (44), with $b$ replaced by $b(1-\varphi_{c}/\varphi)$, namely
@@
$s=\frac{\mathfrak{r}-1}{2}|k|-\mathcal{M}_{c}^{\prime}k^{2},$ (57)
since $\alpha\to 0$, $\gamma\to D/2$, and $k_{c}\to(\mathfrak{r}-1)/2D$; the instability onset is then determined by $\mathfrak{r}=1^{+}$, so that $D/2=\mathcal{M}_{c}^{\prime}$. The dispersion relation (57) then leads to the classical Michelson–Sivashinsky equation
$f_{t}+\tfrac{1}{2}f_{x}^{2}-\mathcal{M}_{c}^{\prime}f_{xx}-\frac{\mathfrak{r}-1}{2}\mathbb{I}(f)=0.$ (58)

### v2

# Flame dynamics and Markstein numbers in Hele-Shaw cells and porous media under Darcy’s law

@@ 3.4 Evolution of unstable planar flames in the presence of a uniform flow
$b$ $=\left[\frac{2\mathfrak{r}}{1+\mathfrak{m}}+\frac{1-\mathfrak{m}}{1+\mathfrak{m}}(\mathcal{V}-1)+\frac{\mathfrak{r}-1}{1+\mathfrak{m}}\frac{\mathfrak{m}}{\mathfrak{r}}\mathrm{g}\right]\mathcal{M}_{t}+\frac{\mathfrak{r}+\mathfrak{m}}{1+\mathfrak{m}}\mathrm{g}\mathcal{M}_{g},$ (38)
@@
  4. 4.
    A key observation is that the dispersion relation no longer involves the curvature Markstein number $\mathcal{M}_{c}$ as defined in formula (10). This is because the factor $(1-\mathbf{v}\cdot\mathbf{n})\nabla\cdot\mathbf{n}$ vanishes then to linear order in the stability analysis, leaving only $\mathcal{M}_{t}$ and $\mathcal{M}_{g}$ contributions. Another aspect worth emphasising is that viscosity variations ($\mu\neq 1$) are ultimately responsible for the dependence of the dispersion relation on the uniform flow magnitude $\mathcal{V}$; indeed, when $\mu=1$, we have $\mathcal{M}_{c}=\mathcal{M}_{t}$ and $\mathcal{V}$ disappears from equations (36) to (39). Consequently, when $\mu=1$, the linear stability of the flame becomes insensitive to the presence of a uniform flow, recovering the usual behaviour common to unconfined flames.
@@
$b$ $=\left(\frac{2\mathfrak{r}}{1+\mathfrak{m}}-\frac{1-\mathfrak{m}}{1+\mathfrak{m}}+\frac{\mathfrak{r}-1}{1+\mathfrak{m}}\frac{\mathfrak{m}}{\mathfrak{r}}\mathrm{g}\right)\mathcal{M}_{t}+\frac{\mathfrak{r}+\mathfrak{m}}{1+\mathfrak{m}}\mathrm{g}\mathcal{M}_{g}.$ (41)
@@
$b$ $=\left(\frac{2\mathfrak{r}}{1+\mathfrak{m}}+\frac{\mathfrak{r}-1}{1+\mathfrak{m}}\frac{\mathfrak{m}}{\mathfrak{r}}\mathrm{g}\right)\mathcal{M}_{t}+\frac{\mathfrak{r}+\mathfrak{m}}{1+\mathfrak{m}}\mathrm{g}\mathcal{M}_{g}.$ (43)
@@ 3.5 Weakly nonlinear regime: the Michelson–Sivashinsky equation for Darcy flows
Near the instability threshold ($a\gtrsim 0$), the unstable wavenumbers scale as $k=O(a)$ when $a\to 0^{+}$, and the growth rate becomes $s\approx a|k|-bk^{2}=O(a^{2})$, with $b=(\mathcal{M}_{t}+\mathrm{g}\mathcal{M}_{g})(\mathfrak{r}+\mathfrak{m})/(1+\mathfrak{m})$ remaining $O(1)$. This scaling mirrors that of the classical Darrieus–Landau instability, but the coefficients now incorporate Darcy-specific contributions. Following the approach of Sivashinsky ashinsky1988nonlinear, michelson1977nonlinear, one obtains a Michelson–Sivashinsky (MS) equation for the flame shape deviation $f(x,t)$ from its unperturbed planar value $y=(\mathcal{V}-1)t$, namely
@@ 4 Flame stability based on Euler–Darcy’s description
$c_{1}$ $=2k+2\mathfrak{r}\mathcal{M}_{t}k^{2}+\varphi(1+ck),$ (47)
$c_{0}$ $=-a\varphi k+\big(b\varphi+1-\mathfrak{r}\big)k^{2}+(3\mathfrak{r}-1)\mathcal{M}_{t}k^{3}$ (48)
@@
$\varphi=\frac{12\textit{Pr}(1+\mathfrak{m})}{\varepsilon^{2}\mathfrak{m}},\quad\quad\varepsilon=\frac{\sqrt{12\kappa_{u}}}{\delta_{L}}.$ (49)
@@
The onset criterion for instability of the planar flame is now given as a function of $\varepsilon$ (or $\varphi$). Following daou2025hydrodynamic, we assume $\mathcal{M}_{t}>0$ (so that $c_{1},c_{2}>0$). The flame is unstable when $a>a_{c}$, with
$a_{c}$ $=\begin{cases}0&\text{if }\varphi>\varphi_{c},\\ \displaystyle\frac{-b(\mathfrak{r}-1)}{4(3\mathfrak{r}-1)\mathcal{M}_{t}}\,\frac{\varphi_{c}}{\varphi}\left(1-\frac{\varphi}{\varphi_{c}}\right)^{\!2}&\text{if }\varphi<\varphi_{c}\end{cases}$ (50)
@@
$k_{c}$ $=\begin{cases}0&\text{if }\varphi>\varphi_{c},\\ \displaystyle\frac{\mathfrak{r}-1}{2(3\mathfrak{r}-1)\mathcal{M}_{t}}\left(1-\frac{\varphi}{\varphi_{c}}\right)&\text{if }\varphi<\varphi_{c}.\end{cases}$ (51)
The first criterion, $a>a_{c}=0$, applies under strong confinement (large $\varphi$) and is predicted by Darcy’s law; it remains valid up to the critical value $\varphi_{c}$. For $\varphi<\varphi_{c}$, the wavenumber at onset is non-zero, with $k_{c}\to(\mathfrak{r}-1)/2(3\mathfrak{r}-1)\mathcal{M}_{t}$ as $\varphi/\varphi_{c}\to 0$, corresponding to unconfined flames.
@@
where $\alpha=\varphi k_{c}/c_{1}(k_{c})$ and $\gamma=(3\mathfrak{r}-1)\mathcal{M}_{t}k_{c}/c_{1}(k_{c})$ are positive numbers. For $\varphi>\varphi_{c}$, the flame dynamics is governed by a Michelson–Sivashinsky-type equation (44), with $b$ replaced by $b(1-\varphi_{c}/\varphi)$, namely
@@
$s=\frac{\mathfrak{r}-1}{2}|k|-\mathcal{M}_{t}k^{2},$ (56)
since $\alpha\to 0$, $\gamma\to\mathcal{M}_{t}$, and $k_{c}\to(\mathfrak{r}-1)/4\mathcal{M}_{t}$ with the instability onset now identified by $\mathfrak{r}=1^{+}$. The dispersion relation (56) then leads to the classical Michelson–Sivashinsky equation
$f_{t}+\tfrac{1}{2}f_{x}^{2}-\mathcal{M}_{t}f_{xx}-\frac{\mathfrak{r}-1}{2}\mathbb{I}(f)=0.$ (57)